### v1

# Measurement, and Mitigation

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, yet they also exhibit memorization of their training data. This phenomenon raises critical questions about model behavior, privacy risks, and the boundary between learning and memorization. Addressing these concerns, this paper synthesizes recent studies and investigates the landscape of memorization, the factors influencing it, and methods for its detection and mitigation. We explore key drivers, including training data duplication, training dynamics, and fine-tuning procedures that influence data memorization. In addition, we examine methodologies such as prefix-based extraction, membership inference, and adversarial prompting, assessing their effectiveness in detecting and measuring memorized content. Beyond technical analysis, we also explore the broader implications of memorization, including the legal and ethical implications. Finally, we discuss mitigation strategies, including data cleaning, differential privacy, and post-training unlearning, while highlighting open challenges in balancing the minimization of harmful memorization with utility. This paper provides a comprehensive overview of the current state of research on LLM memorization across technical, privacy, and performance dimensions, identifying critical directions for future work.
1 Introduction
Throughout the past few years, we have observed great strides in the capabilities of LLMs driven by changes in model architecture, training methodologies, and computational resources (Radford et al., 2018; Brown et al., 2020; Chowdhery et al., 2023; Naveed et al., 2023; Touvron et al., 2023; Wei et al., 2023). These advances have significantly improved natural language understanding, generation, and reasoning abilities, enabling these models to perform increasingly complex tasks across diverse domains (Alowais et al., 2023; Khurana et al., 2023; Minaee et al., 2025). However, in addition to their strong capabilities, LLMs continue to manifest critical limitations when evaluated on criteria such as privacy and security (Chang et al., 2024; Yao et al., 2024a; Das et al., 2025).
With the advent of data scale used to train LLMs, the risks of privacy leakages through memorization have dramatically increased, as models trained on massive datasets containing potentially sensitive information may inadvertently reproduce verbatim passages from training data when prompted with specific triggers, compromising individual privacy and confidentiality without explicit consent (Dwivedi et al., 2023; Smith et al., 2023; Abdali et al., 2024). Thus, data memorization emerges as a critical vulnerability in LLMs, presenting significant privacy risks, including potential exposure of personally identifiable information (PII), copyright violations, and unauthorized reproduction of sensitive content (Carlini et al., 2021; Cremer et al., 2022; Clusmann et al., 2023; Karamolegkou et al., 2023; Song et al., 2023; Freeman et al., 2024; Hua et al., 2024; Morris et al., 2025).
This investigation explores memorization mechanisms in LLMs, examining contributing factors, detection methodologies, measurement approaches, and mitigation techniques. Unlike previous surveys on security and privacy (Siau and Wang, 2020; Fui-Hoon Nah et al., 2023; Guo et al., 2023; Hadi et al., 2023; Karabacak and Margetis, 2023; Min et al., 2023; Zhao et al., 2023; Zhu et al., 2023; Chang et al., 2024; Liu et al., 2024a; Myers et al., 2024; Raiaan et al., 2024; Stahl and Eke, 2024; Zhao et al., 2024), we present a focused examination of memorization phenomena in LLMs, dissecting the fundamental components contributing to and mitigating unintended data retention (Bender et al., 2021; Zhao et al., 2022a; Hartmann et al., 2023; Satvaty et al., 2024). This paper gives an in-depth review of the existing research on LLM memorization and its outcomes. Opposed to the survey by Satvaty et al. (2024), our analysis examines the technical underpinnings, evaluation methods, and privacy implications of memorization while providing concrete research directions. Figure 1 provides a visual taxonomy that maps and categorizes memorization, detection & mitigation of memorization, and corresponding risks.
Figure 1: Taxonomy of understanding memorization in LLMs.
2 Defining memorization
Agnostic to LLMs, memorization occurs when a model can reproduce specific training sequences when provided with appropriate contextual inputs (Carlini et al., 2023; Cooper et al., 2023; Schwarzschild et al., 2024). Currently, several definitions are used to quantify memorization in LLMs, including memorization defined by string matching (exact or approximate) or by loss differences.
2.1 Exact memorization
Verbatim memorization refers to a model’s exact reproduction of training data, often arising due to the high duplication of specific examples or overfitting (Carlini et al., 2021). This phenomenon occurs when an LLM, instead of generating novel responses or synthesizing information, simply retrieves and reproduces specific text snippets word-for-word from the data it was exposed to.
Perfect memorization describes a model that has exactly recorded its training data, assigning a probability only to inputs it has previously seen (Kandpal et al., 2022). When such a model generates an output, the process is mathematically equivalent to randomly selecting examples from its training dataset with a uniform probability.
Eidetic memorization (Carlini et al., 2021) occurs when a prompt $p$ causes the model to reproduce a verbatim string $s$ from training data. A stricter variant, $k$-eidetic memorization (Carlini et al., 2021), occurs when $s$ appears in no more than $k$ training examples, yet can still be extracted given the right prompt.
Discoverable memorization. A string $s$ is discoverably memorized (Carlini et al., 2023; Nasr et al., 2023) when, given a training example composed of prefix $p$ and suffix $s$, a LLM exactly reproduces the continuation $s$ that followed $p$ in the training data.
2.2 Approximate memorization
Approximate/Paraphrased memorization (Ippolito et al., 2023) is when LLMs generate outputs similar to training data in content, structure, or phrasing without exact replication. This differs from exact copying through variations, paraphrasing, or partial overlaps. The authors detect memorization by calculating edit distance between a generation and a target string, normalized by length, and choosing a threshold that determines when a sequence is deemed an approximate match.
2.3 Prompt-based memorization
Extractable memorization occurs when, without access to the training data, there exists a constructable prompt that invokes the model to generate an exact example from its training set (Nasr et al., 2023).
$k$-extractable memorization represents a stricter form of extractable memorization. A suffix is $k$-extractable (Biderman et al., 2023). when, given only its corresponding $k$-token prefix, the model reproduces the entire suffix verbatim. Unlike verbatim memorization, which manifests as direct replication, $k$-extractable memorization captures a model’s ability to retrieve and complete specific training sequences when prompted with only partial context.
($n,p$)-discoverable extraction (Hayes et al., 2025) formalizes the likelihood of retrieving a memorized string via repeated sampling. A string is ($n,p$)-discoverable if it appears in at least one of $n$ completions with probability $\geq p$.
2.4 Influence-based memorization
Counterfactual memorization (Feldman and Zhang, 2020; Zhang et al., 2023) quantifies how much a model’s predictions change based on whether a model is trained on a particular training example. For example, this can be instantiated as how much the loss changes on a datapoint when seen by a model during training versus not. Pappu et al. (2024) instantiate this definition by classifying datapoints as memorized when both a model trained on the datapoint achieves an approximate match under prefix-based decoding and a model not trained on that datapoint does not achieve an approximate match. As it is computationally prohibitive to retrain a model per datapoint-exlusion, Zhang et al. (2023) and Pappu et al. (2024) train a smaller number of models that exclude entire subsets of data used to test for memorization.
3 Factors influencing memorization
With the increasing scale and capability of LLMs, a range of training and inference-time factors have been identified as influencing memorization behavior. This section systematically analyzes the primary variables that govern memorization.
Model parameter size. Model size strongly correlates with increased memorization in LLMs, with larger models demonstrating both greater capacity to retain training data and increased vulnerability to extraction attacks. Carlini et al. (2021) first introduced the relationship between model size and memorization, observing that memorization scales log-linearly with model size. Subsequent research has consistently reinforced this finding across multiple dimensions of memorization. For instance, Tirumala et al. (2022) observed that larger LLMs not only memorize more content but do so more rapidly during training, a phenomenon that cannot be fully explained by conventional overfitting or hyper-parameter tuning. This enhanced memorization capability directly impacts privacy concerns, as Li et al. (2024) demonstrated that extraction attacks become significantly more effective against larger LLMs due to their advanced memorization capabilities. The scaling effect has been further corroborated by (Lee et al., 2022; Tirumala et al., 2022; Carlini et al., 2023; Freeman et al., 2024; Kiyomaru et al., 2024; Li et al., 2024; Hayes et al., 2025; Morris et al., 2025).
Training data duplication. Duplicate examples in training data skew LLMs toward overrepresented content, diminishing output diversity and increasing verbatim reproduction of training material (Carlini et al., 2021). As demonstrated in Lee et al. (2022), de-duplication substantially reduces memorization, with models trained on original data exhibiting a tenfold increase in memorized token generation compared to those trained on datasets, where exact substrings and near-duplicates were removed. Exploring further, Kandpal et al. (2022) identified a superlinear relationship between training data duplication and memorization in LLMs, where rarely duplicated training samples are seldom memorized. However, models generate duplicated sequences at rates significantly lower than perfect memorization would predict.
Sequence length. Longer sequences increase memorization in LLMs. Carlini et al. (2023) found that memorization increases logarithmically with sequence length, with the verbatim reproduction probability rising by orders of magnitude as sequences extend from 50 to 950 tokens. Similarly, extraction methodologies, including continuous soft prompting and dynamic soft prompting techniques, demonstrate a consistent pattern wherein the volume of extracted data grows proportionally with increasing prefix token sizes (Wang et al., 2024).
Tokenization. Models trained with larger Byte Pair Encoding (BPE) vocabularies were found to memorize significantly more sequences from their training data. Kharitonov et al. (2021) studied how different vocabulary sizes, produced via BPE, affect a transformer LLMs’ tendency to memorize sequences. Memorization was particularly strong for named entities, URLs, and uncommon phrases, often becoming single tokens under larger BPE settings.
Sampling Methods. Sampling methods significantly impact the extraction of memorized content from LLMs, with stochastic approaches consistently outperforming greedy decoding in revealing memorized training data. While Carlini et al. (2021) initially established that greedy decoding can reveal memorized sequences but misses others due to its low diversity, subsequent research has demonstrated the superior effectiveness of other sampling methods. Yu et al. (2023) showed that optimizing sampling parameters like top-k, nucleus sampling, and temperature can substantially increase memorized data extraction, in some cases doubling previous baselines. This finding was reinforced by Tiwari and Suh (2025), who discovered that randomized decoding nearly doubles leakage risk compared to greedy decoding, contradicting earlier assumptions. Hayes et al. (2025) provided a theoretical foundation for these observations by introducing a probabilistic framework that demonstrates how repeated sampling with varied decoding parameters can expose memorization hidden under greedy approaches. Given that no single decoding method minimizes leakage across all scenarios, Tiwari and Suh (2025) emphasized the importance of assessing memorization under diverse sampling strategies.
While data duplication has been broadly associated with increased memorization, it remains uncertain whether this effect systematically varies across different data modalities (e.g., code, prose, etc.). Additionally, although inference hyper-parameters, such as decoding methods, are known to impact memorization, existing techniques do not reliably prevent the extraction of memorized content. This underscores the need for decoding strategies explicitly designed to minimize memorization. In contrast, the influence of training hyper-parameters on memorization remains poorly understood, particularly with regards to the role of overfitting through learning rate and regularization mechanisms. Moreover, the temporal dynamics of memorization in the presence of data duplication remain underexplored. Specifically, it is unclear how the timing of when duplicated examples are encountered during training affects memorization likelihood.
Open Questions 1. How do data duplication and data type affect memorization extent? 2. What is required to design a decoding method that minimizes leakage? 3. What factors cause larger LLMs to memorize more? 4. How can we distinguish useful generalization from memorization in LLMs? 5. How do training hyperparameters influence memorization? 6. Does the temporal order of data presentation during training affect memorization likelihood?
4 Memorization at different stages
Although memorization has often been attributed to overfitting, this section explores how it is more deeply shaped by pre-training dynamics and fine-tuning strategies in LLMs.
Pre-training dynamics introduce systematic biases that influence which training examples are retained. Under standard non-deterministic training regimes, characterized by data shuffling, dropout, and stochastic optimization, models tend to forget examples seen early in training unless they are revisited, as shown in Jagielski et al. (2023), Kiyomaru et al. (2024) and Leybzon and Kervadec (2024). They hypothesize that this forgetting is not simply due to limited exposure but rather the result of parameter drift, whereby updates driven by later data overwrite earlier representations. Consequently, memorization is biased toward examples encountered in the latter stages of training.
This effect becomes even more pronounced as models approach the end of training. Huang et al. (2024) showed that later-stage checkpoints are more susceptible to memorizing even rare or out-of-distribution content, likely reflecting increased model capacity and representational flexibility at those stages. Extending this perspective, Biderman et al. (2023) demonstrated that memorization follows predictable scaling laws. As model size and training duration increase, specific sequences predictably transition from unmemorized to memorized. This behavior reinforces that memorization is not an artifact of overfitting but an emergent property of scale and optimization dynamics.
Supervised fine-tuning influences memorization behavior, shaping both the extent and nature of information retention and exposure. Mireshghallah et al. (2022) provided a comparative analysis of fine-tuning approaches, demonstrating that head-only fine-tuning presents the highest risk of memorization, likely due to overfitting. In contrast, adapter-based fine-tuning, when constraining parameter updates (Zeng et al., 2024), reduces memorization using parameter-efficient methods, demonstrating fine-tuning can be effective for limiting privacy risks. They link differences in memorization from models fine-tuned for tasks such as summarization vs. question answering to the attention dynamics of each task, with narrower attention patterns correlating with higher memorization. From an adversarial perspective, Chen et al. (2024) demonstrated that fine-tuning can also be leveraged to extract memorized pretraining data. Using the Janus Interface, they show that targeted fine-tuning amplifies a model’s capacity to leak sensitive information, thus framing fine-tuning as a mechanism that can reactivate latent memorization vulnerabilities. Together, these studies underscore that fine-tuning critically shapes memorization patterns.
Reinforcement learning as post-training. Post-training paradigms such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024), and Reinforcement Learning from Internal Feedback (RLIF) (Zhao et al., 2025) leverage reinforcement learning to train LLMs from various feedback sources. Limited research exists on memorization propagation through reinforcement learning stages. In the context of code generation, Pappu et al. (2024) found that data memorized during fine-tuning persists with high frequency in post-RLHF models, while finding minimal evidence that reward model data or reinforcement learning data becomes memorized. Critical questions remain regarding the predictability of memorization persistence between fine-tuning and reinforcement learning based on data attributes.
Distillation is another common technique used in modern ML pipelines (Hinton et al., 2015; Zhao et al., 2022b), where a small ‘student’ model is trained on data produced by a larger ‘teacher’ model (typically logits). Chaudhari et al. (2025) show that bias injected adversarially into teacher models can be amplified in student models via distillation. This naturally suggests that memorization can propagate between teacher and student models, but this has not been formally analyzed.
Characterizing memorization across training phases requires understanding the mechanisms by which information is encoded, retained, or discarded. Although memorization tends to intensify in later training stages, the degree to which it can be attenuated or reversed remains an open question. With the increasing prevalence of fine-tuned models, it is also unclear how established memorization scaling laws extend to domain-adapted settings. Moreover, the relative contributions of training dynamics versus fine-tuning objectives to overall memorization behavior are not yet well understood.
Open Questions 1. How can we reduce parameter drift to retain early training knowledge? 2. Do memorization scaling laws apply to domain-specific fine-tuned models? 3. How can we quantify and attribute memorization to pre-training vs. fine-tuning? 4. Are there identifiable factors that predict what type of data will persist as memorized between pre-training and post-training? 5. How does memorization persist (or not) from teacher to student models during distillation?
5 Detecting memorization
A wide range of techniques has been developed to detect memorization in LLMs. This section categorizes and examines these methods in depth to offer a comprehensive understanding of how memorization is identified.
Divergence attack. Nasr et al. (2023) proposed a divergence attack, a prompt-based extraction method that circumvents alignment defenses in instruction-tuned LLMs, coaxing the model toward reverting to pre-aligned behavior. This divergence significantly increases the likelihood of emitting memorized training data, achieving up to 150$\times$ more verbatim sequences compared to typical user queries. Nasr et al. (2023) hypothesized that this vulnerability arises because prompts induce decoding analogous to an end-of-text token during pretraining, a context in which LLMs are known to favor high-likelihood and often memorized continuations. The attack exposes persistent memorization that may have remained latent.
Prefix-based data extraction attack works by querying a model with the initial segment of a memorized sequence and observing whether it completes the rest (verbatim or approximate, depending on the specific definition used). First proposed by Carlini et al. (2021), they demonstrated this technique on GPT-2, successfully recovering hundreds of memorized examples by prompting the model with partial strings from the training set. Carlini et al. (2023) showed that longer prefixes increased the likelihood of verbatim completions by reducing ambiguity and aligning the model more strongly with memorized content. Building on Carlini et al. (2023), Li et al. (2024) introduced the LLM-PBE toolkit, a benchmark for evaluating privacy risks via prefix-based extraction. Li et al. (2024)’s findings reinforced that structured prefixes, such as email headers or document beginnings, are potent at eliciting memorized sequences.
An adversarial variant of this attack was explored by Kassem et al. (2024), who used LLMs to generate candidate prefixes likely to elicit private data from a target model. Kassem et al. (2024) observed that instruction-tuned models can leak pretraining data even when the prompts diverge from the original training distribution, suggesting that memorization in LLMs may be more pervasive than what is revealed by prefix attacks. An extension of prefix attacks is soft prompting, which is explored in further detail below.
Membership inference attack (MIA) (Shokri et al., 2017) has emerged as a key diagnostic tool for detecting memorization in LLMs, offering insight into whether specific data points may have been memorized during training. To circumvent the computational overhead of training shadow models, recent work has focused on black-box strategies that exploit the model’s output probabilities or perplexity for input sequences (Duan et al., 2024).
A number of MIA techniques rely on loss- or likelihood-based metrics. Yeom et al. (2018) used the raw loss on a target input, assuming seen examples have lower loss. To mitigate confounding effects from input difficulty, reference-based calibration (Carlini et al., 2021) subtracts the loss from a separate reference model, aiming to isolate differences due to training exposure. Another method used in Carlini et al. (2021) is zlib entropy, which normalizes loss by the zlib-compressed size of the input sequence, using compression as a proxy for complexity. Moving beyond single-point estimates, the neighborhood attack (Mattern et al., 2023) examines local changes in the loss landscape. By generating nearby perturbations of the target input and comparing their average loss, it identifies instances where the target appears memorized. Similarly, the min-$k$prob method (Shi et al., 2023) focuses on the subset of tokens with the highest loss—under the assumption that even the most uncertain tokens in a memorized sequence will still be predicted with confidence.
Yet, despite their utility in highlighting patterns of memorization, MIAs fall short when used as per-instance indicators of training data inclusion. As Zhang et al. (2025) and Duan et al. (2024) argued, MIAs lack a well-calibrated null model since one cannot feasibly train an identical model without the target input. This makes it impossible to meaningfully estimate false positive rates, undermining the statistical soundness of individual predictions. This limitation is underscored by the critique in Maini and Suri (2024), which challenges the claims made in Zhang et al. (2024). The latter introduces the PatentMIA benchmark and a divergence-based calibration method for MIAs, reporting improved reliability in detecting whether a sequence appeared in pretraining. However, Maini and Suri (2024) showed these gains are likely artifacts of distributional shifts introduced by the benchmark itself, which MIAs can exploit without truly resolving the underlying calibration problem. In light of these concerns, Zhang et al. (2025) recommended that MIAs be re-framed as tools for aggregate-level privacy auditing rather than as evidence of individual training data exposure. For stronger guarantees, they advocate for alternatives, such as data extraction attacks or canary-based MIAs, which offer more direct insights into memorization and leakage.
Soft prompting techniques have emerged as powerful tools, enabling both the extraction and suppression of memorized content. By leveraging learned embeddings pre-pended to model inputs, continuous soft prompts can explicitly influence what a model reveals from its training data. For instance, Ozdayi et al. (2023) demonstrated that fixed-length continuous prompts could be trained to either amplify or diminish memorization, as measured by the extraction rates of secret sequences. Their results showed that attack prompts increased memorization leakage by up to 9.3%, while suppress prompts decreased extraction by up to 97.7%. Building on Ozdayi et al. (2023), Kim et al. (2023) introduced ProPILE, a privacy auditing framework that uses soft prompt tuning in a white-box setting to extract PII memorized by LLMs. Their learned prompts were transferable across models, suggesting consistent memorization patterns that can be systematically exploited.
Dynamic soft prompting techniques extend the static nature of continuous prompts by conditioning on the input context. Wang et al. (2024) proposed prefix-conditioned prompting, where a prompt generator produces tailored soft prompts based on the input prefix. This dynamic approach allows LLMs to adapt to subtle input variations and surface context-dependent memorized completions. Wang et al. (2024) revealed that dynamic prompts significantly improved the discoverable memorization rate, surpassing both static prompt and no-prompt baselines. These results underscore the role of context-sensitive prompting in uncovering latent memorized content and highlight the broader utility of soft prompting in analyzing memorization in LLMs.
Reasoning. Much of recent literature has been focused on training Large Reasoning Models (LRMs), models trained to reason before producing a final answer, as opposed to immediately producing an answer in their first token. To date, most memorization literature has been focused on quantifying memorization of specific datapoints by way of measuring whether the model can reproduce the training example. As the field moves towards reasoning models, a definition of what it means for a model to memorize a reasoning pattern may be useful for guiding reasoning research and ensuring reasoning models generalize. Huang et al. (2025) takes a step towards this by measuring whether models trained to reason on math are able to solve slightly perturbed versions of problems in their training set. The authors find that models fail to solve problems that constitute ‘hard’ perturbations, where the problem looks superficially similar to a problem in the training set but requires a different reasoning strategy. Creating benchmarks to measure memorization of such superficial reasoning patterns remains an open and useful direction for LRM research.
From the memorization detection strategies explored above, except for the divergence attack, adversaries either imbue or infer/know the training data intended for extraction. Therefore, it is critical to develop methods to reliably detect memorized sequences in language models without requiring access to the original training data. Furthermore, these detection methods all examine direct memorization from training data. With more powerful LLMs, the boundary between memorization and generalization is fuzzier, propelling the need to understand if memorization can be detected from patterns using reasoning.
Open Questions 1. How can we reliably detect memorization without training data access? 2. Can LLMs infer training data from context rather than direct memorization? 3. How can we detect memorization of implicit patterns and sensitive relationships? 4. How can we build benchmarks to measure whether reasoning models memorize superficial reasoning patterns or learn general reasoning principles?
6 Mitigating memorization
Mitigation techniques for memorization span various approaches, from data handling to model training and auditing. This section outlines the main categories, explores key strategies within each, and highlights their differences in effectiveness and trade-offs.
6.1 Training-time interventions
Data cleaning plays a crucial role in mitigating memorization by limiting an LLM’s exposure to overrepresented sequences. As demonstrated by Carlini et al. (2021), Lee et al. (2022), and Kandpal et al. (2022), de-duplication minimizes the over-representation of repeated sequences. Removing these duplicates serves as a form of implicit regularization, reducing the likelihood that the LLM overfits.
From an optimization perspective, LLMs can minimize loss on rare training examples by memorizing. This makes PII-scrubbing effective: by removing these sequences during training, memorization is minimized (Carlini et al., 2021; Jagielski et al., 2023).
Differential privacy (DP) provides robust protection against membership inference and extraction attacks, even in the presence of adversaries with auxiliary knowledge (Dwork, 2006; Dwork et al., 2014; Shokri et al., 2017; Carlini et al., 2021). Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016) guarantees privacy by computing per-example gradients, clipping them to a fixed norm to bound individual influence, and injecting calibrated Gaussian noise into the aggregated updates. A privacy accountant tracks cumulative privacy loss to enforce the global ($\epsilon$, $\delta$) budget.
Li et al. (2021) demonstrated that LLMs can act as strong DP learners. Their evaluation reveals that pretrained LLMs are highly resilient to the noise introduced by DP-SGD, particularly when fine-tuned on top of general-purpose representations. DP-trained models approach the performance of non-private baselines, while offering provable protection against memorization. Empirical validation through canary insertion and MIAs confirm the effectiveness of DP in preventing leakage of sensitive sequences. To further reduce memorization, Zhao et al. (2022a) introduced Confidentially Redacted Training (CRT), a method that combines de-duplication and redaction with DP-SGD to train LLMs while avoiding the retention of sensitive content. CRT builds on DP concepts to introduce randomized training interventions that prevent unintended memorization. The authors show that CRT, when combined with DP-SGD, provides strong privacy protections without compromising model performance, as indicated by competitive perplexity scores.
However, the use of DP in LLMs often presents trade-offs between privacy, utility, and computational efficiency. Strict privacy budgets often lead to performance drops, and significant computational overhead makes it challenging to scale to LLMs. To address these challenges, recent work has examined parameter-efficient fine-tuning (PEFT) strategies for training under DP. PEFT approaches, such as LoRA (Hu et al., 2022), allow finetuning of models with addition of few parameters relative to base models via low-rank decompositions of additional trainable parameters. It is hypothesized that finetuning PEFT models with DP mitigates losses of utility due to requiring less noise for privatization due to training with fewer parameters (Liu et al., 2024b; Yu et al., 2022, 2021). Ma et al. (2024) explored the integration of adapter-based fine-tuning methods into the DP framework. Their findings reveal that DP PEFT reduces memorization and achieves comparable or superior downstream task performance relative to full-model DP-SGD, under strict privacy budgets. Ma et al. (2024) hypothesized that the limited parameter updates in PEFT may concentrate DP noise in a narrow subset of the model, potentially weakening privacy protection. While most DP techniques for LLMs target record-level privacy, many applications require user-level guarantees. Chua et al. (2024) addressed this gap by proposing several techniques to enforce DP at user-level. To evaluate effectiveness, Chua et al. (2024) used canary insertion attacks, embedding unique sequences into user data to test for memorization. Results show that user-level DP reduces memorization, with much lower canary extraction rates than record-level DP.
DP in LLMs continue to present trade-offs between privacy, utility, and computational efficiency. Strict privacy budgets often lead to performance drops, and significant computational overhead makes it challenging to scale to LLMs. Thus, the adoption of DP for mitigating memorization in LLMs remains limited.
6.2 Post-training-time interventions
Machine unlearning aims to remove the influence of certain training examples so that the model’s behavior is as if those examples were never seen (Cooper et al., 2024; Liu, 2024).  Yao et al. (2024b) evaluates various model-editing unlearning strategies (e.g., gradient ascent (Golatkar et al., 2020; Jia et al., 2023; Jang et al., 2022), negative re-labeling (Golatkar et al., 2020), adversarial sampling (Cha et al., 2024)) demonstrating that approximate unlearning methods can be over $10^{5}\times$ more computationally efficient than retraining a model from scratch. However, unlike DP, there is no formal guarantee, thereby leaving a risk that memorization persists.
ParaPO, introduced by Chen et al. (2025), is a post-training strategy to decrease memorization of data memorized in the pretraining corpus. ParaPO decreases unintended memorization by first assessing which data from the pretraining corpus has been memorized (by searching for sequences that, when a prefix is used to prompt an LM, an LM decodes a near exact match to the suffix of the sequence), and then creating preference pairs by using a separate LLM to summarize the memorized datapoint. The model is then post-trained via DPO (Rafailov et al., 2024) on pairs of (memorized sequence, summarization), with the summarization marked as the preferred response. Chen et al. (2025) show that this approach decreases unintended memorization while preserving verbatim recall of desired sequences (e.g. direct quotations) However, the authors note that this approach slightly decreases utility on math, knowledge and reasoning benchmarks.
Machine unlearning and ParaPO constitute post-training interventions for reducing memorization. ParaPO results in slight degradation of utility, raising the question of what extent memorization is required for utility and generalization. Additionally, a future research direction may be to what extent this tradeoff can be tuned via incorporation of “memorization” reward models that are optimized in conjunction with utility-oriented reward models, in the context of RLHF or RLVR.
6.3 Inference-time interventions
MemFree decoding is a strategy introduced by Ippolito et al. (2023) to filter out memorized sequences during generation. This approach contrasts with decoding by introducing a privacy filter into the generation loop applied post hoc at generation and acts as a wrapper around any pre-trained model. Ippolito et al. (2023) leveraged a bloom filter to represent the set of all n-grams in the training set to identify memorized content in real time. However, near-identical n-grams are undetectable as minor modifications to a token sequence evade detection. Another limitation is that access to the LLM’s n-gram training data is required.
Activation steering addresses memorization in LLMs by manipulating internal activations during inference (Turner et al., 2023). By injecting targeted perturbations into an LLM’s hidden states, activation steering can suppress or redirect the generation of memorized sequences while preserving overall model reasoning capabilities (Suri et al., 2025). Suri et al. (2025) demonstrated that using a sparse autoencoder to identify activation patterns linked to memorized passages enables the construction of steering directions that reduced memorization by up to 60%, with minimal performance degradation. However, they note that steering effectiveness depends heavily on precise layer selection and steering strength.
Understanding where memorization arises in LLMs is crucial to designing effective steering interventions. Stoehr et al. (2024) conducted a detailed mechanistic analysis of paragraph-length memorized sequences and found that memorization is often localized to specific LLM components. They showed that fine-tuning only a subset of the high-gradient weights was sufficient to erase memorized passages. Additionally, they identified a single attention head in an early layer that reliably activates in response to rare token combinations seen during training, triggering verbatim recall.
Building on localization, Chang et al. (2023) evaluated multiple methods for localizing memorized content by introducing two diagnostic benchmarks: injection, where known memorized sequences are inserted via fine-tuning a small set of weights, and deletion, which tests whether removing specific neurons erases a naturally memorized output. They compared several techniques, including activation-based heuristics (Geva et al., 2022), integrated gradients (Dai et al., 2021), and pruning-based methods like Slimming and Hard Concrete (Chang et al., 2023). Their results showed that pruning-based approaches were most effective: for example, Hard Concrete was able to identify fewer than 0.5% of neurons whose removal led to a  60% drop in memorization accuracy. Both Chang et al. (2023) and Stoehr et al. (2024) highlighted a key challenge: the neurons involved in one memory often contribute to others, complicating memory-specific interventions due to the risk of collateral forgetting.
By leveraging insights from model localization and sparse representation learning, such techniques can target memorized content with increasing precision, offering a flexible and minimally invasive tool for minimizing memorization.
While data cleaning should be a fundamental part of every LLM training pipeline to reduce memorization, other mitigation strategies either face practicality and scalability trade-offs or lack formal guarantees of effectiveness. Memorization mitigation is typically approached from two perspectives: (1) addressing all training data, and (2) targeting specific subsets of information intended to be forgotten. Given that LLMs often indiscriminately memorize data, mitigation efforts should prioritize reducing the retention of sensitive or harmful content without compromising knowledge essential for factual recall and reasoning. From the first perspective, ensuring robust privacy protection requires evaluating the viability of scaling up differentially private (DP) training and understanding the optimal trade-off between utility and cost. Additionally, as PEFT methods grow in popularity, there is a clear need to refine DP strategies to suit these approaches and mitigate some of the overhead associated with full-model DP training. For targeted mitigation, activation steering shows promise in localizing specific knowledge representations. However, further work is needed to formalize layer selection and parameter tuning to improve our understanding and control of memorization in LLMs.
6.4 Connections to other undesirable behaviors
It remains open whether mitigating memorization also mitigates other undesirable behaviors of LLMs, such as hallucination, racial bias, and toxic content. For instance, does reducing memorization reduce hallucination on concepts related to memorized data? Is it possible that reducing the memorization of datapoints that include harmful racial stereotypes reduces the rates at which large models produce biased or toxic content? The authors to date have not seen cross-cutting analysis examining whether reducing memorization can reduce other harmful behaviors in-tandem.
Open Questions 1. How can LLMs be designed or trained for selective memorization? 2. Shall we scale up DP training, and at what cost and utility trade-off? 3. How can DP improve parameter-efficient fine-tuning to reduce memorization? 4. How can we optimize activation steering for memorization removal? 5. How can post-training methods reduce memorization while preserving utility? 6. Can memorization mitigation methods be incorporated online during post-training, for instance, by using memorization-detecting reward models during RLHF? 7. When does memorization contribute to versus harm utility/generalization? 8. Does mitigating memorization affect other undesirable behaviors, such as hallucination or generating toxic stereotypes?
7 Privacy & legal risks of memorization
Memorization in LLMs poses serious security and privacy risks: memorized text can be directly leaked, undermining data confidentiality (Smith et al., 2023). We discuss three major impact areas caused by memorization: personal data leakage, exposure of copyrighted or proprietary content, and broader legal and public consequences.
Personal data leakage. As explored in Li et al. (2024), Carlini et al. (2021), Jagielski et al. (2023), and Zhou et al. (2024), a significant risk for LLMs is leaking sensitive personal information. In Panda et al. (2024), researchers demonstrated a targeted neural phishing attack that achieved up to a 50% success rate in tricking a model into revealing PII during fine-tuning when injecting poisons during pretraining. Even when training data is intended to remain private, malicious attackers with the correct model access level can siphon memorized secrets (Panda et al., 2024). If an LLM reveals any PII, this could result in identity theft or other harms. Furthermore, for industries such as healthcare or customer service, where customers expect confidentiality, this would place model operators in the crosshairs of current privacy regulations.
Copyright or proprietary content. Beyond PII, memorization in LLMs may also cause the reproduction of copyrighted or proprietary material such as books, code, etc. This turns the LLM into a potential conduit for the unauthorized distribution of protected content. As observed by Carlini et al. (2023), models may produce verbatim memorization from their training data. While we can only explicitly confirm what training data is used for open-source LLMs (Gao et al., 2020; Soldaini et al., 2024), there do exist copyrighted books in EleutherAI’s PILE dataset, specifically, copyrighted books in its Books3 section. This is observed in Henderson et al. (2023) and Zhang et al. (2022), where open models can regurgitate the first few pages of the Harry Potter books and “Oh the places you’ll go!” by Dr. Seuss verbatim, raising questions regarding copyright.
Legal landscape. There are several ongoing cases that will shape the legal landscape of LLMs. The New York Times v. Microsoft Corp. lawsuit (NYT-MC, 2023) represents a pivotal legal challenge to generative AI and copyright, accompanied by similar cases, including Chabon v. OpenAI (Chabon-OpenAI, 2023) where authors claiming copyright infringement in model training, and Doe v. GitHub (DOE-GitHub, 2022) where plaintiffs allege verbatim code reproduction in violation of the DMCA. The Times claimed that OpenAI’s model not only trained on its copyrighted articles without proper license but also that ChatGPT could regurgitate substantial passages of those articles on request. In NYT-MC (2023), the publishers argue this behavior amounts to mass copyright infringement since the model’s outputs mimic and even compete with the Times’ content. While researchers such as  Freeman et al. (2024) may provide their analysis regarding the veracity of lawsuit (NYT-MC, 2023)’s copyright infringement claims, such high-profile legal action underscores how memorization is no longer only a technical but a business and public policy problem.
Open Questions 1. Can LLMs be trained to avoid memorizing copyrighted content? 2. How can memorization metrics inform legal analysis (fair use, copyright)? 3. What technical threshold defines legally significant memorization?
8 Conclusion
Memorization in LLMs represents a double-edged sword in the advancement of AI systems. LLMs demonstrate significant capacity to retain and reproduce training data, with larger models exhibiting higher memorization rates of verbatim content, PII, and copyrighted material. This phenomenon creates a fundamental tension between model utility and privacy/legal concerns that must be addressed.
Detection methods for memorization, like prefix-based extraction and membership inference attacks, remain limited, requiring new approaches that work without training data access and standardized auditing tools for comprehensive evaluation. While mitigation strategies, including de-duplication, scrubbing, differential privacy, and model unlearning, show potential, we must better understand their impact on model utility. Ongoing legal challenges highlight the need for balanced frameworks that preserve beneficial knowledge retention while preventing information leakage and establishing appropriate memorization boundaries.
By advancing our understanding in these areas, we can work toward building more powerful and trustworthy AI systems that maintain high performance while respecting privacy and intellectual property concerns. This balance will be essential as LLMs continue to be integrated into increasingly sensitive applications.